# Addressing divergent representations from causal interventions on neural networks

### v2

@@ 4.2 Pernicious divergence via off-manifold activation
We call a hidden pathway any unit, vector direction, or subcircuit that is inactive on the natural support of representations but becomes active and influences behavior under an intervention. Hidden pathways are compatible with benign divergences (Sec. 4.1), yet they undermine claims about native mechanisms, and they can prime dormant behavioral changes (Sec. 4.2.2).
@@
4.2.1 Mean-difference patching activates hidden pathways
@@
If we project $\hat{h}^{\ell}_{B\rightarrow A}$ onto $\mathrm{conv}(S_{A})$ (or equivalently onto a local PCA subspace of $S_{A}$), this ReLU state change disappears, and the decision boundary is no longer crossed—confirming that the original effect was driven by an off-manifold activation rather than a genuine within-manifold causal mechanism.
General condition.
Let $u_{k}^{\top}$ denote row $k$ of $W$. If $u_{k}^{\top}h+b_{k}\leq 0$ for all $h\in S_{A}$ but $u_{k}^{\top}(h_{B}+\delta_{B\rightarrow A})+b_{k}>0$ for some $h_{B}$, then $\delta_{B\rightarrow A}$ activates a hidden pathway.
@@ 4.2.2 Dormant behavioral changes
Here, the argmax index of $\hat{y}=[\hat{y}_{1},\hat{y}_{2},\hat{y}_{3}]^{\top}$ corresponds to class predictions A, B, and C.
@@
Dormant behavioral changes highlight that behaviorally “safe” interventions can still introduce hidden context dependencies. Detecting them would require evaluating across all possible contexts, which is infeasible in practice. Therefore, interventions should (1) report any newly active units relative to native ReLU patterns, (2) project perturbed representations onto local manifolds, and (3) test for context-sensitive output flips across nuisance dimensions.
@@
Hidden pathways arise when off-manifold interventions activate neurons that never fire for native representations. Such pathways can alter causal conclusions even when behavior appears unchanged. Manifold-preserving projections and ReLU-pattern audits provide practical safeguards against these pernicious forms of divergence.
@@ 5 How might we avoid divergent representations?
We have thus far demonstrated that divergent representations are empirically common and that they can theoretically be harmless or pernicious depending on the functional landscape of the target model and the causal claims. In this section, we consider the question of how such divergences might be avoided. There are existing methods that constrain counterfactual features (in our case, intervened representations) directly to the natural manifold (verma2024counterfactualexplanationreview). However, we seek a method that generates principled interventions on intermediate representations that are constrained to be innocuous. In pursuit of this method, we modify and apply the Counterfactual Latent (CL) loss introduced in grant2025mas. We apply this in synthetic settings as a demonstration of how to reduce pernicious divergence by minimizing all divergence along causally relevant dimensions.
5.1 Synthetic Model Setting
We simulate an intermediate layer of an NN by constructing a synthetic dataset of $h$ vectors with known feature dimensions and labels $y$ that we use to train a Multi-Layer Perceptron (MLP). The dataset consists of noisy samples around a set of grid points from two feature dimensions with correlational structure. Specifically, we define a base set of points given by the Cartesian product of two values along the $x_{1}$-axis, $\{-1,1\}$, and five evenly spaced values along the $x_{2}$-axis, $\{0,1,2,3,4\}$, resulting in ten unique base coordinates, each corresponding to 1 of 10 classes. We add noise and covariance to these feature dimensions and concatenate $n$ extra noise dimensions resulting in simulated vectors $h\in\mathbb{R}^{2+n}$, where $n=64$ unless otherwise stated. The feature dimensions of these synthetic vectors are shown as the natural distributions in Figures 1(b) and 3. We then train a small MLP on these representations to predict the class labels using a standard cross entropy loss. After training, we perform DAS with varying intensities of the behavioral and CL loss term. See Appendix A.2 for further details on the dataset construction, MLP training, and DAS trainings.
5.2 Adapted Counterfactual Latent Loss
@@
To encourage intervened representations to be more similar to the native distribution of NN representations, we build upon the CL auxiliary loss from grant2025mas. This auxiliary objective relies on counterfactual latent (CL) vectors as vector objectives. CL vectors are defined as vectors that encode the same causal variable value(s) that would exist in the intervened vector, $\hat{h}$, assuming the interchange intervention was successful. We can obtain CL vectors by searching through a pre-recorded set of natural $h$ vectors for situations and behaviors that are consistent with the values of the causal abstraction to which we are aligning. See Figure 3(a) for a visualization. As an example, assume that we have a CA with variables $\text{var}_{u}$, $\text{var}_{w}$, and $\text{var}_{\textit{extra}}$, and following a causal intervention we expect $\hat{h}$ to have a value of $u$ for variable $\text{var}_{u}$ and $w$ for variable $\text{var}_{w}$. The CL vector $h_{\textit{CL}}$ in this example can be obtained from averaging over a set of $m$ natural representations, $h_{\textit{CL}}=\frac{1}{m}\sum_{i=1}^{m}h^{(x_{i})}_{\textit{CL}}$, where each $h^{(x_{i})}_{\textit{CL}}$ has the same variable values: $\text{var}_{u}=u$ and $\text{var}_{w}=w$ labeled by the CA. Note that grant2025mas constructed $h_{\textit{CL}}$ using $m=1$, whereas we use $m$ equal to all valid $h^{(x_{i})}_{\textit{CL}}$ in the dataset.
@@
The CL auxiliary loss $\mathcal{L}_{\textit{CL}}$ introduced in grant2025mas is composed of the mean of an L2 and a cosine distance minimization objective using CL vectors as labels. In this work, we modify this loss by applying it to individual causal subspaces. This allows us to construct $h_{\textit{CL}}^{\text{var}_{i}}$ vectors specific to a single causal variable $\text{var}_{i}$. Using notation defined in Section 2.2, the $\mathcal{L}_{\textit{CL}}$ for a single training sample is defined as follows:
[…]
$z^{\text{var}_{i}}_{\textit{CL}}$ $=$ $D_{\text{var}_{i}}\mathcal{A}(h_{\textit{CL}})$ (8)
$\mathcal{L}_{\textit{CL}}$ $=$ $\sum_{i=1}^{n}\mathcal{L}^{\text{var}_{i}}_{\textit{CL}}=\sum_{i=1}^{n}\Big(\frac{1}{2}\|\hat{z}^{\text{var}_{i}}-z^{\text{var}_{i}}_{\textit{CL}}\|^{2}_{2}-\frac{1}{2}\frac{\hat{z}^{\text{var}_{i}}\cdot z^{\text{var}_{i}}_{\textit{CL}}}{\|\hat{z}^{\text{var}_{i}}\|_{2}\,\|z^{\text{var}_{i}}_{\textit{CL}}\|_{2}}\Big)$ (9)
@@
The CL loss is combined with the DAS loss into a single loss term using $\epsilon$ as a tunable hyperparameter: $\mathcal{L}_{total}=\epsilon\mathcal{L}_{\textit{CL}}+\mathcal{L}_{\textit{DAS}}$. The loss is computed as the mean over batches of samples and optimized using gradient descent (Appendix A.2).
@@
Figure 3: The CL loss reduces representational divergence and improves out-of-distribution generalization. (a) Diagram of the CL loss. The rectangles represent model latent vectors. $x_{1}$ and $x_{2}$ indicate the latent state’s non-noisy values of the two synthetic, causal dimensions shown in panels (b) and (c), labeled by the causal abstraction. We patch the $x_{2}$ value from the source vector into the target using DAS. The modified counterfactual latent (CL) loss vector is defined as the average over all naturally occurring latent vectors that possess the same variable values as the post-intervention latent vector. (b) Each panel shows the two causal feature dimensions of hidden states (represented by each dot) from a synthetic dataset. Colors indicate the class of the hidden state; dark dots indicate that the state is from the natural distribution while lighter dots indicate that the state is from the intervened distribution. The leftmost panel shows results from DAS trained using behavior only (Equation 3), the rightmost panel shows DAS trained on only the CL loss (Equation 9). We can see that the CL loss reduces the divergence along the feature dimensions.
5.3 The CL loss alleviates divergence and can improve generalization
@@
Figures 3(b) and  3(c) provide a qualitative comparison of intervened and native representations for interventions using a trained DAS rotation matrix. Each dot in the figures shows the values of the feature dimensions for a single representation. The native states are displayed in darker colors and the intervened in lighter. Each hue indicates the ground truth class of the state. We can see a tightening of the intervened representations with increasing intensity of the CL loss. Quantitatively, the DAS loss produces EMD values along the feature dimensions of $0.043\pm 0.002$ whereas the CL loss produces $0.024\pm 0.001$ with IIAs of $0.981\pm 0.003$ and $0.977\pm 0.004$ respectively.
@@
We hypothesized that reduced divergence could improve IIA when transferring the DAS rotation matrix to OOD settings. To test this, we partitioned the synthetic task into a Dense and a Sparse group, named for the relative spacing of 4 distinct classes within each group. See Appendix A.2.4 for further details on the partitions. We then performed DAS on each of the partitions and evaluated the performance on the held-out partition to produce the results in Figure3(d). The CL loss performs better than the behavioral loss in these OOD settings. We provide further exploration of the CL loss IIA and EMD across a range of hyperparameters in Appendix A.2.5.
@@ 6 Discussion and limitations
In this work we demonstrated that a variety of common causal interventions can produce representations that diverge from a target model’s original distribution of latent representations. We then showed that although such divergences can have benign effects for many causal claims, they can also activate hidden pathways and trigger dormant behaviors that can perniciously affect other claims. As a step towards mitigating this issue, we provided a broad-stroke solution by directly minimizing the divergence of intervened activity along causal dimensions, mitigating both pernicious and harmless forms of divergence.
A remaining gap in our work is the failure to produce a reliable method for classifying pernicious forms of divergent representations. Additionally, the CL loss that we have presented is confined to a narrow set of simplistic settings and is not specific to pernicious divergence. We look forward to exploring ways to classify and mitigate pernicious divergence through self-supervised means in future work.
@@ A.1.2 Measuring Divergence
To quantify distributional shifts between original and intervened representations, we approximated the Earth Mover’s Distance (EMD) (villani2009wassersteinEMD) using the Sinkhorn loss from the GeomLoss python package with a $p=2$ and $\text{blur}=0.05$ (cuturi2013sinkhorn). Let $\mathcal{H}=\{h_{i}\}_{i=1}^{N}$ denote a set of original representations and $\hat{\mathcal{H}}=\{\hat{h}_{i}\}_{i=1}^{N}$ their intervened counterparts. We computed
$\text{EMD}(\mathcal{H},\hat{\mathcal{H}})=\min_{\gamma\in\Pi(\mu,\nu)}\sum_{i,j}\gamma_{ij}\,\|h_{i}-\hat{h}_{j}\|_{2},$ (12)
where $\mu$ and $\nu$ are the empirical distributions over $\mathcal{H}$ and $\hat{\mathcal{H}}$, and $\Pi(\mu,\nu)$ denotes the set of couplings with marginals $\mu$ and $\nu$.
Baseline correction.
To isolate divergence due specifically to intervention, we subtracted a baseline EMD, computed by comparing the original set $\mathcal{H}$ to a separately sampled set $\mathcal{H}^{\prime}$:
$\text{Divergence}=\text{EMD}(\mathcal{H},\hat{\mathcal{H}})-\text{EMD}(\mathcal{H},\mathcal{H}^{\prime}).$ (13)

### v3

@@ 4.2 Pernicious divergence via off-manifold activation
Pernicious cases of representational divergence involve the concept of hidden pathways, which refer to any unit, vector direction, or subcircuit that is inactive on the natural support of representations for a given context but becomes active and influences behavior under an intervention. Although hidden pathways can be compatible with harmless divergences (Section 4.1), they can also undermine claims about natural mechanisms and can prime dormant behavioral changes discussed in Section 4.2.2.
@@
4.2.1 Mean-difference patching can activate hidden pathways
@@
If we project $\hat{h}^{\ell}_{B\rightarrow A}$ onto $\mathrm{conv}(S_{A})$ (or equivalently onto a local PCA subspace of $S_{A}$), this ReLU state change disappears, and the decision boundary is no longer crossed—confirming that the original effect was driven by an off-manifold activation rather than a genuine within-manifold causal mechanism. This experiment does not support claims about the natural mechanisms of the model.
@@ 4.2.2 Dormant behavioral changes
Here, the first argmax index of $\hat{y}=[\hat{y}_{1},\hat{y}_{2},\hat{y}_{3}]^{\top}$ corresponds to class predictions A, B, and C.
@@
Dormant behavioral changes highlight that behaviorally “safe” interventions can still introduce hidden context dependencies. Detecting them would require evaluating across all possible contexts, which is infeasible in practice. Therefore, causal intervention based experiments should ideally (1) report any introduced representational divergence outside of the null-space and (2) test causal interventions for context-sensitivity.
@@
Hidden pathways arise when off-manifold representations activate computations that never occur for natural representations. Such pathways can alter causal conclusions even when behavior appears unchanged depending on the claims. Manifold-preserving projections and ReLU-pattern audits provide practical safeguards against these pernicious forms of divergence.
@@ 5 How might we avoid divergent representations?
We have thus far shown that divergent representations are empirically common and that they can be harmless or harmful depending on multiple factors. We now consider the question of how such divergences might be avoided. Some existing methods solve this by constraining counterfactual features (in our case, intervened representations) directly to the natural manifold (verma2024counterfactualexplanationreview). However, we seek a method that generates principled interventions that are constrained to be innocuous. In this pursuit, we first apply the Counterfactual Latent (CL) loss from grant2025mas to the DAS experiments from wu2024alpacadas and find that we can reduce intervened divergence while preserving accuracy on a Llama based LLM. We then introduce a modified CL loss that only targets causal dimensions, and we show that it can improve OOD intervention accuracy.
5.1 Applying the Counterfactual Latent Loss to Boundless DAS
@@
To encourage intervened NN representations to be more similar to the native distribution, we first apply the CL auxiliary loss from grant2025mas to the Boundless DAS setting in wu2024alpacadas. This auxiliary training objective relies on counterfactual latent (CL) vectors as vector objectives. CL vectors are defined as vectors that encode the same causal variable value(s) that would exist in the intervened vector, $\hat{h}$, assuming the interchange intervention was successful. We can obtain CL vectors from sets of natural $h$ vectors from situations and behaviors that are consistent with the values of the CA to which we are aligning. See Figure 3 for a visualization. As an example, assume that we have a CA with causal variables $\text{var}_{u}$, $\text{var}_{w}$, and $\text{var}_{\textit{extra}}$, and following a causal intervention we expect $\hat{h}$ to have a value of $u$ for variable $\text{var}_{u}$ and $w$ for variable $\text{var}_{w}$. A CL vector $h_{\textit{CL}}$ for this example can be obtained by averaging over a set of $m$ natural representations, $h_{\textit{CL}}=\frac{1}{m}\sum_{i=1}^{m}h^{(x_{i})}_{\textit{CL}}$, where each $h^{(x_{i})}_{\textit{CL}}$ has the same variable values: $\text{var}_{u}=u$ and $\text{var}_{w}=w$ (as labeled by the CA).
@@
The CL auxiliary loss $\mathcal{L}_{\textit{CL}}$ introduced in grant2025mas is composed of the mean of an L2 and a cosine distance objective using CL vectors as labels. Using notation defined in Section 2.2, $\mathcal{L}_{\textit{CL}}$ for a single training sample is defined as follows:
$\mathcal{L}_{\textit{CL}}(\hat{h},h_{CL})=\frac{1}{2}\|\hat{h}-h_{\textit{CL}}\|^{2}_{2}-\frac{1}{2}\frac{\hat{h}\cdot h_{\textit{CL}}}{\|\hat{h}\|_{2}\,\|h_{\textit{CL}}\|_{2}}$ (7)
@@
$\mathcal{L}_{\textit{CL}}$ is combined with the DAS behavioral loss $\mathcal{L}_{\textit{DAS}}$ into a single loss term using $\epsilon$ as a tunable hyperparameter: $\mathcal{L}_{total}=\epsilon\mathcal{L}_{\textit{CL}}+\mathcal{L}_{\textit{DAS}}$. The loss is computed as the mean over batches of samples and optimized using gradient descent (Appendix A.4).
Results: We applied the CL loss to the Boundless DAS notebook from wu2024pyvene which reproduces the main result from wu2024alpacadas (see Appendix A.3). Figure 3A provides qualitative visualizations of the decreased divergence when applying the CL loss, and Figure 3B shows both IIA and EMD as a function of increasing the CL weight $\epsilon$. For small values of $\epsilon$, the IIA is maintained (and potentially even improved) while EMD decreases, demonstrating that the CL auxiliary loss can directly reduce divergence in practical interpretability settings without sacrificing the interpretability method.
5.2 Modified CL loss improves OOD interventions in synthetic model settings
We next modify the CL loss to work independently of the DAS behavioral loss and show that it improves OOD intervention performance. We simulate an intermediate layer of an NN by constructing a synthetic dataset of $h$ vectors with known feature dimensions and labels $y$ that we use to train a Multi-Layer Perceptron (MLP). The dataset consists of noisy samples around a set of grid points from two feature dimensions with correlational structure. Specifically, we define a set of features as the Cartesian product of two values along the $x_{1}$-axis $\{-1,1\}$ and five values along the $x_{2}$-axis $\{0,1,2,3,4\}$, resulting in ten unique coordinates that each correspond to one of ten classes. We add noise and covariance to these feature dimensions and concatenate $n$ extra noise dimensions, resulting in simulated vectors $h\in\mathbb{R}^{2+n}$ where $n=16$ unless otherwise stated. The feature dimensions of these vectors are shown as the natural distributions in Figures 1(b) and 3. We then train a small MLP on these representations to predict the class labels using a standard cross entropy loss. After training, we perform DAS analyses with either the behavioral loss or the CL loss independently. See Appendix A.4 for further details on dataset construction, MLP training, and DAS training.
We modify the CL loss by applying it to individual causal subspaces only (as discovered through the DAS training). This allows us to construct $h_{\textit{CL}}^{\text{var}_{i}}$ vectors specific to a single causal variable $\text{var}_{i}$. The modified $\mathcal{L}^{\prime}_{\textit{CL}}$ for a single training sample is defined as follows:
$\hat{h}^{\text{var}_{i}}=\mathcal{A}^{-1}(D_{\text{var}_{i}}\mathcal{A}(\hat{h})),\quad h^{\text{var}_{i}}_{\textit{CL}}=\text{stopgrad}(\mathcal{A}^{-1}(D_{\text{var}_{i}}\mathcal{A}(h_{\textit{CL}})))$ (8)
$\mathcal{L}^{\prime}_{\textit{CL}}=\sum_{i=1}^{n}\mathcal{L}^{\text{var}_{i}}_{\textit{CL}}=\sum_{i=1}^{n}\Big(\frac{1}{2}\|\hat{h}^{\text{var}_{i}}-h^{\text{var}_{i}}_{\textit{CL}}\|^{2}_{2}-\frac{1}{2}\frac{\hat{h}^{\text{var}_{i}}\cdot h^{\text{var}_{i}}_{\textit{CL}}}{\|\hat{h}^{\text{var}_{i}}\|_{2}\,\|h^{\text{var}_{i}}_{\textit{CL}}\|_{2}}\Big)$ (9)
@@ 5.2 Modified CL loss improves OOD interventions in synthetic model settings
Figure 3: The CL loss reduces representational divergence and can improve out-of-distribution generalization. (a) PCA of natural (orange) and intervened (blue) representations in the Boundless DAS setting presented in wu2024alpacadas for two CL loss weightings with the same final IIA. (b) IIA (orange) and divergence (purple) of intervened representations from Section 5.1 as a function of CL loss weight ($\epsilon$). (c) Diagram of CL loss; rectangles are model representations and $x_{1}$ and $x_{2}$ are deterministic values of the representations along the two synthetic causal dimensions shown in panels (d) and (e). We patch the $x_{2}$ value from source to target using DAS and define the CL representation as the average of all natural representations that possess the same variable values as the post-intervention representation. (d) and (e) two causal feature dimensions of representations from a synthetic dataset consisting of ten classes (colors), with both natural (dark) and intervened (light) representations shown. (d) shows results from DAS trained using behavior only, (e) shows DAS trained using only the CL loss. (f) performance of alignment matrices trained on one task and evaluated on another that uses the same causal dimensions. CL loss leads to higher OOD performance.
@@
Results: Figures 3D and  3E provide a qualitative comparison of intervened and native representations for interventions using a trained DAS rotation matrix. Each dot in the figures shows the values of the feature dimensions for a single representation. The native states are displayed in darker colors and the intervened in lighter. Each hue indicates the ground truth class of the state. We can see a tightening of the intervened representations using the CL loss. Quantitatively, the DAS loss produces EMD values along the feature dimensions of $0.032\pm 0.003$ whereas the CL loss produces $0.007\pm 0.001$ with IIAs of $0.997\pm 0.001$ and $0.9988\pm 0.0005$ respectively on training/test sets with held-out classes.
@@
We hypothesized that reduced divergence could improve IIA when transferring the DAS rotation matrix to OOD settings. To test this, we partitioned the synthetic task into a dense and a sparse cluster of classes based on the relative spacing of 4 distinct classes within each cluster (see Appendix A.4.4 for partition details). We then performed a separate MLP and DAS training on each partition and evaluated the IIA on the held-out partition to produce Figure 3F. The CL loss performs better than the behavioral loss in these OOD settings. We provide CL loss ablations in Appendix A.4.5.
@@ 6 Discussion and limitations
In this work we demonstrated that a variety of common causal interventions can produce representations that diverge from a target model’s natural distribution of latent representations. We then showed that although this can have benign effects for many causal claims, it can also activate hidden pathways and trigger dormant behaviors that can perniciously affect other claims. As a step towards mitigating this issue, we provided a broad-stroke solution by directly minimizing the divergence of intervened activity along causal dimensions, mitigating both pernicious and harmless forms of divergence.
A remaining gap in our work is the failure to produce a principled, reliable method for classifying harmful divergence. Additionally, the modified CL loss that we have presented is confined to a narrow set of simplistic settings and is not specific to pernicious divergence. We look forward to exploring ways to classify and mitigate pernicious divergence through self-supervised means in future work.
@@ 7 Acknowledgments
8 LLM Usage Statement
We used ChatGPT to elicit editing suggestions for clarity, as well as to improve notational consistency and clarity in the hidden pathways and dormant behavioral changes formalizations in Sections 4.1 and 4.2. We also used ChatGPT to provide an initial layout of the proof offered in Appendix A.2 showing that axis-aligned hyperrectangles are the only manifold shape that do not have divergent source-target vector pairs in coordinate patching settings.
@@ A.1.2 Measuring Divergence
For each intervened sample, there exists a corresponding ground truth sample that the intervention is meant to approximate. In the case of the mean difference experiments, these ground samples consist of the naturally occurring entity or attribute in the position which the $\delta_{MD}$ is meant to approximate. For the SAEs, each reconstructed vector has a corresponding encoded vector. For DAS, the ground truth vectors are equivalent to CL vectors.
Earth Mover’s Distance: To quantify distributional differences between original and intervened representations, we approximated the Earth Mover’s Distance (EMD) (villani2009wassersteinEMD) including all vector dimensions using the Sinkhorn loss from the GeomLoss python package with a $p=2$ and $\text{blur}=0.05$ (cuturi2013sinkhorn). Let $\mathcal{H}=\{h_{i}\}_{i=1}^{N}$ denote a set of original representations and $\hat{\mathcal{H}}=\{\hat{h}_{i}\}_{i=1}^{N}$ their intervened counterparts. We computed
$\text{EMD}(\mathcal{H},\hat{\mathcal{H}})=\min_{\gamma\in\Pi(\mu,\nu)}\frac{1}{N}\sum_{i,j}\gamma_{ij}\,\|h_{i}-\hat{h}_{j}\|_{2},$ (12)
where $N$ is the number of samples in $\mathcal{H}$, $\mu$ and $\nu$ are the empirical distributions over $\mathcal{H}$ and $\hat{\mathcal{H}}$, and $\Pi(\mu,\nu)$ denotes the set of couplings with marginals $\mu$ and $\nu$.
Baseline divergence.
To ensure that we did not introduce bias from the sampling procedure in the baseline comparison, we use the corresponding ground truth vectors for the intervened vectors when comparing to the baseline divergence. Formally, let $\mathcal{H}^{\prime}$ be the set of ground truth natural vectors corresponding to the set $\hat{\mathcal{H}}$:
$\text{Divergence}_{\text{Baseline}}=\text{EMD}(\mathcal{H},\mathcal{H}^{\prime}).$ (13)
Local PCA Distance:
For each reference point $x_{i}$, we identified its $k$ nearest neighbors $\mathcal{N}_{k}(x_{i})$ in Euclidean space and computed the local tangent subspace via PCA. Let $U_{d}(x_{i})\in\mathbb{R}^{D\times d}$ denote the top $d$ principal components explaining at least 95% of the local variance. The projection matrix onto this tangent space is $P_{i}=U_{d}U_{d}^{\top}$.
For a query point $v$, the Local PCA Distance is the orthogonal residual between $v$ and its projection onto the local tangent space at the nearest reference sample:
$D_{\mathrm{PCA}}(v)=\min_{x_{i}}\left\|(I-P_{i})\,[\,v-x_{i}\,]\right\|_{2}.$
Small values indicate that $v$ lies close to the locally linear approximation of the manifold, whereas large residuals reflect departures orthogonal to the manifold surface.
Local Linear Reconstruction Error:
We computed an error inspired by Locally Linear Embedding (LLE). Given the same neighborhood $\mathcal{N}_{k}(v)$ of $k$ reference points $X_{k}=\{x_{1},\dots,x_{k}\}$, we found reconstruction weights $w=(w_{1},\dots,w_{k})$ that best express $v$ as a convex combination of its neighbors:
$\min_{w}\ \left\|v-\sum_{j=1}^{k}w_{j}x_{j}\right\|_{2}^{2}\qquad\text{s.t.}\qquad\sum_{j}w_{j}=1.$
A small Tikhonov regularizer $\lambda I$ was added to the local covariance for numerical stability. The Local Linear Reconstruction Error is the residual norm at the optimum:
$D_{\mathrm{LLR}}(v)=\left\|v-X_{k}w^{\ast}\right\|_{2}.$
This metric measures how well $v$ can be expressed as a locally linear interpolation of nearby manifold points; poor reconstruction (large $D_{\mathrm{LLR}}$) implies off-manifold position or local curvature mismatch.
Kernel Density (KDE) Density Score:
We estimated a nonparametric probability density function over the reference set using Gaussian kernel density estimation:
$\hat{p}(x)=\frac{1}{nh^{D}}\sum_{i=1}^{n}\exp\!\left[-\frac{\|x-x_{i}\|_{2}^{2}}{2h^{2}}\right],$
where $h$ is the kernel bandwidth determined by Silverman’s rule of thumb or cross-validation. The KDE Density Score for a query point $v$ is its log-density under this model,
$S_{\mathrm{KDE}}(v)=\log\hat{p}(v),$
which inversely reflects off-manifold distance: lower log-density corresponds to less typical or out-of-distribution points. To express results on a comparable scale, we report the negative log-density (i.e., $S_{\mathrm{KDE}}$), so that larger values consistently indicate greater deviation from the manifold.